Using Cayley — Hamilton theorem, evaluate A^(7) if A=[ 3 2
-1. 0 ]
1
Expert's answer
2020-12-21T18:59:47-0500

The characteristics equation is "|A-\lambda I|=0"

"\begin{vmatrix}
 3-\lambda & 2\\
 -1 & \lambda
\end{vmatrix}=0\\
-\lambda(3-\lambda)+2=0\\
-3\lambda+\lambda^2+2=0\\
\lambda^2-3\lambda+2=0"

By Cayley-Hamilton Theorem,

"A^2-3A+2I=0\\
\implies A^2=3A-2I"


"A^7=A^4.A^3\\
A^7=(A^2)^2.A^2.A\\
A^7=(3A-2I) ^2.(3A-2I).A\\
A^7=(9A^2-12A+4I)(3A^2-2A)\\
A^7=(9(3A-2I)-12A+4I)(3(3A-3I)-2A) \\
A^7=(27A-18I-12A+4I)(9A-6I-2A) \\
A^7=(15A+4I)(7A-6I) \\
A^7=(105A^2-188A+84I)\\
A^7=(105(3A-2I)-188A+84I)\\
A^7=315A-210A-188A+84I\\
A^7=127A-126I\\
A^7=127\begin{pmatrix}
 3&2 \\
 -1&0
\end{pmatrix}-126\begin{pmatrix}
 1&0 \\
 0&1
\end{pmatrix}\\
A^7=\begin{pmatrix}
 381&254\\
 -127&0
\end{pmatrix}-\begin{pmatrix}
 126&0 \\
 0&126
\end{pmatrix}\\
A^7=\begin{pmatrix}
255&254 \\
 -127&-126
\end{pmatrix}"